Annual salaries for a large company are approximately normally distributed with a mean of
R50,000 and a standard deviation of R20,000.
a. What salary would an employee need to get in order to be in the lowest 30%?
b. What is the probability of having an above average salary range of between R60000 to
R80000
1
Expert's answer
2019-04-28T15:33:30-0400

If "X\sim N(\mu,\sigma^2)" then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"


"\mu=R50,000,\ \ \ \sigma= R20,000"

a. What salary would an employee need to get in order to be in the lowest 30%? 


"P(X<x^*)=P(Z<z^*)=0.3=>z^*=-0.5244"

"x^*=z^*\sigma+\mu"

"x^*=-0.5244(R20000)+R50000=R39512"


b. What is the probability of having an above average salary range of between R60000 to R80000


"z_1=\dfrac{R60000-R50000}{R20000}=0.5"

"z_2=\dfrac{R80000-R50000}{R20000}=1.5"

"P(R60000<X<R80000)=P(0.5<Z<1.5)="

"=P(Z<1.5)-P(<Z<0.5)=0.93319280-0.69146246="


"=0.24173034"